Circular No. 9020

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 N3 (LULIN) M. R. Combi, University of Michigan: J.-L. Bertaux and E. Quèmerais,

Service d’Aèronomie, Centre National de la Recherche Scientifique and Université de Versailles Saint-Quentin en Yvelines; and J. T. T. Mäkinen, Finnish Meteorological Institute, note that the Solar Wind Anisotropies (SWAN) camera on the Solar and Heliospheric Observatory (SOHO) spacecraft, located at the earth-sun L_{1} Lagrange point, makes daily full-sky images of hydrogen Lyα. SWAN detected comet C/2007 N3 from Jan. 20 to 30. The following water-production rates (given in units of 10^{27} s^{−1}) were determined from the hydrogen Lyα brightness and distribution: Jan. 20.5 UT, 98; Jan. 21.5, 98; Jan. 22, 100; Jan. 23.5, 120; Jan. 24.5, 110; Jan. 25.5, 79; Jan. 26.5, 120; Jan. 27.5, 79; Jan. 28.5, 77; Jan. 29.5, 84; Jan. 30.5,

79. Calibration, background subtraction, and model parameters contribute to a realistic absolute uncertainty of ∼ 30 percent (there are uncertainties from the instrument noise and formal model inversion).

COMETS C/2008 T11, C/2008 T12, AND C/2008 U2–U4 (SOHO) Further to IAUC 9018, additional Kreutz sungrazers have been found

on SOHO website images. C/2008 T11 was small and stellar in appearance (mag ∼ 7–7.5) in C3 images, but diffuse and faint with no tail in C2 images. C/2008 T12 was very small, diffuse, and elongated (mag ∼ 8). C/2008 U2 was small and stellar in appearance (mag ∼ 7). C/2008 U3 was slightly diffuse (mag ∼ 8–8.5). C/2008 U4 was slightly diffuse, appearing small (mag ∼ 7) in C3 images with a short, diffuse tail in C2 images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 T11 Oct. 13.638 13^{h}01.^{m}3 − 9^{o}32′ C3/2 HS 2009-A59
C/2008 T12  14.313 13 11.3 − 9 13 C2  RK 2009-A59
C/2008 U2  16.521 13 19.8 −10 06 C2  EB 2009-B66
C/2008 U3  16.788 13 21.3 −10 09 C2  RM 2009-B66
C/2008 U4  16.863 13 08.1 −11 06 C3/2 RM 2009-B66

COMETS 213P/VAN NESS, 214P/LINEAR, AND 215P/NEAT Comet P/2009 B3 = P/2005 R2 (cf. IAUC 9017) has been assigned

the permanent designation 213P. Comet P/2009 B4 = P/2002 CW_{134} (cf. IAUC 9017) has been assigned the permanent designation 214P. Comet P/2009 B5 = P/2002 O8 (cf. IAUC 9018) has been assigned the permanent designation 215P. See also MPC 64980.

2009 February 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green